# Ergodic Estimates of One-Step Numerical Approximations for Superlinear SODEs

### v1

@@ 3.3. Ergodic Estimates of Numerical Schemes
We apply Theorem 2.3 to this part’s implicit and explicit numerical schemes.
@@
(1) The discrete processes $(Y_{k})_{k\in\mathbb{N}}$ given in Example 3.3 are homogeneous Markov chains and possess unique invariant measures, denoted by $\pi^{M}_{\tau}$ (for MEM) and $\pi^{B}_{\tau}$ (for BEM), respectively [12, 11].
@@
(3.15) $\sup_{\|\varphi\|_{4}\leq 1}|\pi_{\tau}^{M}(\varphi)-\pi(\varphi)|$ $\leq C\tau,$
(3.16) $\sup_{\|\varphi\|_{4}\leq 1}|\pi_{\tau}^{B}(\varphi)-\pi(\varphi)|$ $\leq C\tau.$
@@
Let us first estimate the term $\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(\mathscr{P}(Y_{0}))$ with $Y_{0}=X_{0}\sim\pi_{\tau}^{M}$. As $\pi_{\tau}^{M}$ is an invariant measure of (3.10), we have $\hat{Y}_{\tau}=Y_{1}\sim\pi_{\tau}^{M}$. Thus, $f(\hat{Y}_{\tau})$ and $f(Y_{0})$ share the same law so that $\mathbb{E}^{\pi_{\tau}}[\mathbb{E}[f(\hat{Y}_{\tau}^{x})-f(Y_{0})]|_{x=\mathscr{P}(Y_{0})}]=0$. Then (2.11) implies
@@
(3.19) $\Big|\sum_{i=1}^{6}\mathbb{E}^{\pi_{\tau}}[\mathbb{E}\mathcal{R}_{i}(x,g_{\tau}(x))|_{x=Y_{0}}]\Big|$ $\leq C\|\nabla^{4}f\|(1+\pi_{\tau}^{M}(|\cdot|^{\alpha_{1}}))\tau^{2},$
(3.20) $|\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(\mathscr{P}(Y_{0}))|$ $\leq C\|\nabla f\|\pi_{\tau}^{M}(|\cdot|^{\alpha_{2}})\tau^{2}.$
@@
Finally, we need to remark that $\pi^{M}_{\tau}(|\cdot|^{\max\{\alpha_{1},\alpha_{2}\}})<\infty$, as $\pi_{\tau}^{M}$ possesses all bounded $2p$-moments for any $p\in[1,p^{\star}]\cap\mathbb{N}_{+}$, which follows from the moments estimate of the MEM (3.10) in [12, Theorem 3.2].
@@
Let $Y_{0}\sim\pi_{\tau}$. As $\pi_{\tau}$ is an invariant measure of (2.9), we have $Y_{1}\sim\pi_{\tau}$. Since $\hat{Y}_{0}=g_{\tau}(Y_{0})$ and $\hat{Y}_{\tau}=g_{\tau}(Y_{1})$, $\hat{Y}_{0}$ and $\hat{Y}_{\tau}$ share the same law. Then (2.11) implies $\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(\hat{Y}_{0})=0$.
@@
Finally, we mention that $\pi^{B}_{\tau}(|\cdot|^{2(\gamma+1)})<\infty$, as $\pi_{\tau}^{B}$ possesses all bounded $2p$-moments for any $p\in[1,p^{\star}]\cap\mathbb{N}_{+}$, which follows from the moments estimate of the BEM (3.13) in [11, Lemma 10]. ∎
@@
Using the same arguments as in the proof of Theorem 3.5, we have the following sharp ergodic error estimate for the EM scheme with a unique invariant measure $\pi_{\tau}^{E}$, under the Lipschitz conditions as in [2, 4, 15]:
$\sup_{\|\varphi\|_{\mathcal{C}_{b}^{4}}\leq 1}|\pi_{\tau}^{E}(\varphi)-\pi(\varphi)|\leq C\tau.$

### v2

@@ 3.3. Ergodic Estimates of Numerical Schemes
We now apply Theorem 2.3 to specific implicit and explicit numerical schemes.
@@
(1) The discrete processes $(Y_{k})_{k\in\mathbb{N}}$ given in Example 3.3 are homogeneous Markov chains and possess unique invariant measures, all denoted by $\pi_{\tau}$. They are VV–geometrically ergodic with $V(x):=1+|x|^{2}$, $x\in\mathbb{R}^{d}$; see [11] for the BEM (3.13) and [12] for the TEM (3.11) and the PEM (3.12).
@@
(3.15) $\sup_{\|\varphi\|_{4}\leq 1}|\pi_{\tau}(\varphi)-\pi(\varphi)|$ $\leq C\tau,$
(3.16) $\Big|\frac{1}{N}\sum_{k=0}^{N-1}\mathbb{E}\varphi(Y_{k}^{x})-\pi(\varphi)\Big|$ $\leq C(1+|x|^{2})(\tau+N^{-1}).$
@@
Let us first estimate the term $\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(\mathscr{P}(Y_{0}))$ with $Y_{0}=X_{0}\sim\pi_{\tau}$. As $\pi_{\tau}$ is an invariant measure of (3.10), we have $\hat{Y}_{\tau}=Y_{1}\sim\pi_{\tau}$. Thus, $f(\hat{Y}_{\tau})$ and $f(Y_{0})$ share the same law so that $\mathbb{E}^{\pi_{\tau}}[\mathbb{E}[f(\hat{Y}_{\tau}^{x})-f(Y_{0})]|_{x=\mathscr{P}(Y_{0})}]=0$. Then (2.11) implies
@@
(3.19) $\Big|\sum_{i=1}^{6}\mathbb{E}^{\pi_{\tau}}[\mathbb{E}\mathcal{R}_{i}(x,g_{\tau}(x))|_{x=Y_{0}}]\Big|$ $\leq C\|\nabla^{4}f\|(1+\pi_{\tau}(|\cdot|^{\alpha_{1}}))\tau^{2},$
(3.20) $|\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(\mathscr{P}(Y_{0}))|$ $\leq C\|\nabla f\|\pi_{\tau}(|\cdot|^{\alpha_{2}})\tau^{2}.$
@@
Finally, we need to remark that $\pi_{\tau}(|\cdot|^{\max\{\alpha_{1},\alpha_{2}\}})<\infty$, as $\pi_{\tau}$ possesses all bounded $2p$-moments for any $p\in[1,p^{\star}]\cap\mathbb{N}_{+}$, which follows from the moments estimate of the MEM (3.10) in [12, Theorem 3.2].
@@
Let $Y_{0}\sim\pi_{\tau}$. As $\pi_{\tau}$ is an invariant measure of (2.9), we have $Y_{1}\sim\pi_{\tau}$. Since $\hat{Y}_{0}=g_{\tau}(Y_{0})$ and $\hat{Y}_{\tau}=g_{\tau}(Y_{1})$, $\hat{Y}_{0}$ and $\hat{Y}_{\tau}$ have the same law. Then (2.11) implies $\mathbb{E}^{\pi_{\tau}}\mathcal{A}_{\tau}f_{\varphi}(\hat{Y}_{0})=0$.
@@
Finally, we mention that $\pi_{\tau}(|\cdot|^{2(\gamma+1)})<\infty$, as $\pi_{\tau}$ possesses all bounded $2p$-moments for any $p\in[1,p^{\star}]\cap\mathbb{N}_{+}$, which follows from the moments estimate of the BEM (3.13) in [11, Lemma 10].
(3) The error estimate (3.16) of the ergodic limit follows from the $V$-geometrical ergodicity of these one-step approximations in Remark 3.4 and the error (2.3). ∎
@@
Using the same arguments as in the proof of Theorem 3.5, we obtain the following sharp ergodic error estimates for the EM scheme with a unique invariant measure $\pi_{\tau}$, under Lipschitz conditions as in [2, 4, 16]:
$\sup_{\|\varphi\|_{\mathcal{C}_{b}^{4}}\leq 1}|\pi_{\tau}(\varphi)-\pi(\varphi)|\leq C\tau,\quad\Big|\frac{1}{N}\sum_{k=0}^{N-1}\mathbb{E}\varphi(Y_{k}^{x})-\pi(\varphi)\Big|\leq C(1+|x|^{2})(\tau+N^{-1}).$
@@
4. Numerical Experiments
This section presents comprehensive numerical investigations to validate the ergodicity (via the strong mixing property) and the ergodic estimates for the BEM (3.13) and the TEM (3.11). We consider the following coefficients:
$b(x)=-x-x^{3},\qquad\sigma(x)=\frac{1}{2}\sqrt{x^{2}+1},\quad x\in\mathbb{R}.$
4.1. Verification of Assumptions
We first verify that the chosen coefficients satisfy all assumptions required by our theoretical framework.
For any $x,y\in\mathbb{R}$, direct computation yields
$\langle x-y,b(x)-b(y)\rangle$ $=(x-y)[(-x-x^{3})-(-y-y^{3})]$
$=-(x-y)^{2}-(x-y)^{2}(x^{2}+xy+y^{2})\leq-(x-y)^{2}.$
Moreover, using the Lipschitz continuity of the square-root function,
$|\sigma(x)-\sigma(y)|^{2}=\frac{1}{4}\Big|\sqrt{x^{2}+1}-\sqrt{y^{2}+1}\Big|^{2}\leq\frac{1}{4}|x-y|^{2}.$
Consequently, for any $p^{\star}\geq 2$, we obtain
$\langle x-y,b(x)-b(y)\rangle+\frac{2p^{\star}-1}{2}|\sigma(x)-\sigma(y)|^{2}\leq-\Bigl[1-\frac{2p^{\star}-1}{8}\Bigr]|x-y|^{2}.$
Choosing $L_{1}=1-(2p^{\star}-1)/8>0$ (e.g., $L_{1}=5/8$ when $p^{\star}=2$) ensures the monotonicity condition (3.1) in Assumption 3.1 holds.
Regarding the dissipative condition (3.2), we have
$\langle x,b(x)\rangle+\frac{p^{\star}(2p^{\star}-1)}{2}|\sigma(x)|^{2}$ $=-x^{2}-x^{4}+\frac{p^{\star}(2p^{\star}-1)}{8}(x^{2}+1)$
$\leq L_{2}-L_{3}|x|^{4},$
where one may take $\gamma=3$ and select suitable positive constants $L_{2},L_{3}$. Since $b$ is cubic and $\sigma$ is smooth with derivatives of polynomial growth, Assumption 3.2 is also satisfied. Hence, the coefficient pair $(b,\sigma)$ is admissible under our theoretical setting.
4.2. Strong Mixing Property
To examine the strong mixing property (which implies unique ergodicity) of TEM (3.11) and the BEM (3.13), we fix the step-size $\tau=0.2$ and simulate $N=5000$ steps, corresponding to a final time $T=1000$. Three distinct initial values, $X_{0}=-5,\;5,\;15$, are used. For each scheme, $M=5000$ independent sample paths are generated.
Figure 1. Empirical density functions of TEM (3.11) (left) and spatial BEM (3.13) (right).
The empirical density at the terminal time $T=1000$ is reconstructed via Gaussian kernel density estimation. Figure 1 displays the resulting empirical density functions for the TEM (3.11) (left) and the BEM (3.13) (right), respectively. Despite the disparate initial conditions, the density curves for each scheme exhibit nearly perfect overlap. This visual agreement numerically confirms that both discretizations possess the strong mixing property and, consequently, that their associated Markov chains are uniquely ergodic.
4.3. Weak Convergence Order
We now investigate the convergence order of the ergodic errors (i.e., weak convergence order) in Theorem 3.5 of these two schemes. The test functions are chosen as $\varphi_{1}(x)=\cos(x)$ and $\varphi_{2}(x)=e^{-x^{2}}$, $x\in\mathbb{R}^{d}$, the initial value is fixed as $X_{0}=1$, and the simulation horizon is set to $T=32$. A reference solution is computed using a fine step-size $\tau_{\text{ref}}=2^{-11}$. Comparative simulations are performed with coarser step-sizes $\tau_{i}=2^{-i}$ for $i=3,4,5,6,7$. Expectations are approximated via Monte Carlo averages over $M=20000$ independent sample paths.
Figure 2. Weak convergence orders of TEM (3.11) (left) and spatial BEM (3.13) (right).
Figure 2 presents, in double logarithmic coordinates, the weak errors of the TEM (3.11) and the BEM (3.13) as functions of $\tau$. For TEM, the estimated convergence orders are $1.07$ with $\varphi_{1}(x)=\cos(x)$ and $1.08$ with $\varphi_{2}(x)=e^{-x^{2}}$. For BEM, the corresponding estimates are $1.08$ and $1.02$. These numerical results robustly support the theoretical prediction in Theorem 3.5, which shows that both schemes achieve a first-order ergodic error for SODEs with superlinear coefficients driven by multiplicative noise.